# Automatic Reward Shaping from Confounded Offline Data

### v1

@@ 1 Introduction
Notations. We will consistently use capital letters ($V$) to denote random variables, lowercase letters ($v$) for their values, and cursive $\mathcal{V}$ to denote their domains. Fix indices $i,j\in\mathbb{N}$. We use bold capital letters ($\boldsymbol{V}$) to denote a set of random variables and let $|\boldsymbol{V}|$ denote its cardinality of set $\boldsymbol{V}$. Finally, $\boldsymbol{1}_{\boldsymbol{Z}=\boldsymbol{z}}$ is an indicator function that returns $1$ if event $\boldsymbol{Z}=\boldsymbol{z}$ holds true; otherwise, it returns 00.
@@ 2.1 Potential-Based Reward Shaping
where $r_{h}$ is the reward function in $\mathcal{M}$; $\phi_{h}(\cdot):\mathcal{S}\mapsto\mathbb{R}$ is a real valued potential function and $\phi_{H}(s)=0$. Then,
That is, every optimal policy in $\mathcal{M}$ will also be an optimal policy in $\mathcal{M}^{\prime}$, and vice versa.
@@ 3 Robust Potential Functions from Confounded Offline Data
where ~$\widetilde{\mathcal{R}}_{h}(s,x)=\mathbb{E}[Y_{h}|S_{h}=s,X_{h}=x]$, $\widetilde{\mathcal{T}}_{h}(s,x,s^{\prime})=P_{h}(S_{h+1}=s^{\prime}|S_{h}=s,X_{h}=x)$ and $P_{h}(x|s)=P_{h}(X_{h}=x|S_{h}=s)$ are all empirical estimations from the offline dataset. $b$ is a known upper bound on the reward signal, $Y_{h}\leq b$. Similar to the bounding approach developed in (Zhang & Bareinboim, 2024), we apply the above bounds to the Bellman Optimal Equation (Sutton & Barto, 2018) and arrive at the following Causal Bellman Optimal Equation that calculates the optimal state value optimistically from the confounded offline datasets.
@@ 4 Efficient Online Learning with PBRS
where we use ^$\hat{\mathbb{P}}_{h}^{k_{i}}$ to denote the empirical transition of episode $k_{i}$. That is, for the value of a function mapping from state space to real numbers, $f:\mathcal{S}\mapsto\mathbb{R}$,its value given the next state $s^{\prime}$ in episode $k_{i}$ as input is denoted as, $f(s^{\prime})=(\hat{\mathbb{P}}_{h}^{k_{i}}f)(s,x)$. Similarly, we write the expected value of such function $f$ over the transitions of an online CMDP under reward shaping, $\mathcal{M}^{\prime}$,as $({\mathbb{P}}_{h}f)(s,x)=\mathbb{E}_{\mathcal{M}^{\prime}_{\pi}}[f(s^{\prime})|s,x]$. $\alpha_{t}^{i}$ is the cumulative learning rates defined as $\alpha_{t}^{0}=\Pi_{j=1}^{t}(1-\alpha_{j})$ and $\alpha_{t}^{i}=\alpha_{i}\Pi_{j=i+1}^{t}(1-\alpha_{j})$. Note that we also assume deterministic reward functions for notation wise simplicity (Jin et al., 2018). With the notations above, we bound the difference between the learned Q-value and the optimal Q-value with high probability in both direction.
@@
With probability at least $1-2p$, the difference between the learned Q-value at the beginning of episode $k$ and step $h$ and the optimal Q-value can be bounded as $0\leq Q^{k}_{h}(s,x)-Q^{*}_{h}(s,x)\leq\alpha_{t}^{0}(-Q^{*}_{h}(s,x))+\sum_{i=1}^{t}\alpha^{i}_{t}\left[(\hat{\mathbb{P}}_{h}^{k_{i}}V_{h+1}^{k_{i}}-\hat{\mathbb{P}}_{h}^{k_{i}}V_{h+1}^{*})(s,x)\right]+3b_{t}$ where $b_{t}$ is the UCB bonus as in Algo. 1.
@@
$Q^{k}_{h}(s_{h}^{k},x_{h}^{k})-Q^{*}_{h}(s_{h}^{k},x_{h}^{k})\leq\operatorname{clip}\left[3b_{t}\left|\frac{\Delta(s_{h}^{k},x_{h}^{k})}{2H}\right.\right]$
@@ 5 Experiments
We use tabular state-action representations and the rewards are all deterministic in those environments. There is a step penalty of $-0.1$, $+0.2$ for getting a coin, 00 for reaching the goal, and $-1$ for touching the lava. The episode ends immediately when either a goal/lava grid is reached or the episode length hits the horizon limit. In windy grid worlds, the state transition is determined by the resultant force of both agent’s action and the wind direction if the agent tries to move. For example, if the agent wants to move right when there is a north wind (wind blowing from the north), the agent will end up being in the lower right grid concerning its original location instead of the right-hand side grid. However, the wind direction can only be observed by more capable behavioral agents. In our collected offline datasets, wind direction is thus a hidden confounder forming a CMDP. See App. D for detailed experiment setups and more results.
@@ Acknowledgements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here. One major reason is that the potential-based reward-shaping framework we considered does not alternate the optimal policy of the agent, which is an inherent property of the problem. Our proposed method solely serves as an acceleration for training such agents.
@@ D.1 Experiment Setups
Regarding the environment parameters, for Windy Empty World, the episode length is set to $15$ while the LavaCross series has a horizon of $20$. To compensate for the hard exploration situation, we allow random initial starting states over the whole map walkable area. For training steps, we set a total of $100$K environment steps for Windy Empty World and $20$K for the LavaCross series. We represent the wind distribution in those windy Grid World as a five-tuple meaning the probability mass for (west wind, north wind, east wind, south wind, no wind). We list them as follows,
  1. Windy Empty World: (.1,.1,.1,.1,.6).1.1.1.1.6(.1,.1,.1,.1,.6)( .1 , .1 , .1 , .1 , .6 )
  2. Windy LavaCross (easy): (0,.8,0,0,.2)0.800.2(0,.8,0,0,.2)( 0 , .8 , 0 , 0 , .2 )
  3. Windy Lava Cross (hard): (0,.7,0,0,.3)0.700.3(0,.7,0,0,.3)( 0 , .7 , 0 , 0 , .3 )
  4. Windy Lava Cross (maze): $\begin{cases}(0.,.5,0.,0.,.5)\ \text{if }s[0]=0\\ (.5,0.,0.,0.,.5)\ \text{if }s[1]=7\\ (.15,.15,.15,.15,.4)\ \text{otherwise}\\ \end{cases}$
@@ Appendix G Useful Results for Proofs
In addition, we have ($\sum_{i=1}^{t}\alpha_{t}^{i}=1$ and $\alpha_{t}^{0}=0$ for $t\geq 1$; (2) $\sum_{i=1}^{t}\alpha_{t}^{i}=0$ and $\alpha_{t}^{0}=1$ for $t=0$.
@@ H.2 Causal Bellman Optimal Equation Proof
where ~$\widetilde{\mathcal{R}}_{h}(s,x)=\mathbb{E}[Y_{h}|S_{h}=s,X_{h}=x]$, $\widetilde{\mathcal{T}}_{h}$ is shorthand for $\widetilde{\mathcal{T}}_{h}(s,x,s^{\prime})=P(S_{h+1}=s^{\prime}|S_{h}=s,X_{h}=x)$ and $P(x|s)=P_{h}(X_{h}=x|S_{h}=s)$ are estimated from the offline dataset. And $b$ is a known upper bound on the reward signal, $Y_{h}\leq b$. In this step, we upper bound the next state transition by assuming the best case that for the action not taken with probability $P_{h}(\neg x|s)$, the agent transits with probability $1$ the best possible next state, $\max_{s^{\prime\prime}}V^{*}_{h+1}(s^{\prime\prime})$.
@@
Let the initial difference between the value vectors be $||\overline{V}_{1}-\overline{V}_{2}||_{\infty}^{w}\leq c$ where $c>0$ is a constant. Then we have, for any state $s\in\mathcal{S}_{k(s)}$ where $k(s)$ means the index of the state set that $s$ belongs to,
@@
$||T\overline{V}_{1}-T\overline{V}_{2}||_{\infty}^{w}\leq\gamma c$ (33)
for all $\overline{V}_{1},\overline{V}_{2}$ satisfying $||\overline{V}_{1}-\overline{V}_{2}||\leq c$. Thus, $T$ is a contraction mapping with respect to a weighted max norm.
@@ H.3 Q-UCB with Shaping Regret Analysis Details
where $\mathbb{P}_{h}$ denotes the expected transition w.r.t the true confounded MDP and ^$\hat{\mathbb{P}}_{h}^{k}$ denotes the sample transition experienced by the agent at episode $k$ step $h$.
@@
By Lem. H.3, we have that with probability at least $1-2p$ for all $(s,x,h,k)\in\mathcal{S}\times\mathcal{X}\times[H]\times[K]$,
@@
    By Lem. H.3 and together with the discussion above for the other side, we know (^$(\hat{\mathbb{P}}_{h}^{k_{i}}\phi-{\mathbb{P}}_{h}\phi)(s,x)\geq-\frac{b_{t}}{2}$ with high probability. While by Lem. G.1, $\sum_{i=1}^{t}\alpha_{t}^{i}b_{i}\geq b_{t}$, thus, we have,
@@
We can split trajectories by whether the high probability event in Lem. H.4 happens. When Lem. H.4 happens with at least $1-2p$ probability, Lem. H.7 is also satisfied and we have,
@@
$\sum_{k=1}^{K}Q^{k}_{h}(s_{h}^{k},x_{h}^{k})-Q^{*}_{h}(s_{h}^{k},x_{h}^{k})$ $\leq\sum_{k^{\prime},h^{\prime}\geq h}(1+\frac{1}{H})^{2(h^{\prime}-h)}\operatorname{clip}\left[3b_{N_{h^{\prime}}^{k^{\prime}}}\left|\frac{\Delta(s_{h^{\prime}}^{k^{\prime}},x_{h^{\prime}}^{k^{\prime}})}{2H}\right.\right]$ (89)
$\leq e^{2}\sum_{k,h}\operatorname{clip}\left[3b_{N_{h}^{k}}\left|\frac{\Delta(s_{h}^{k},x_{h}^{k})}{2H}\right.\right].$ (90)
@@
$=e^{2}H\sum_{h=1}^{H}\sum_{(s,x)\in\operatorname{Sub}_{\Delta}}\sum_{i=1}^{N_{h}^{K}(s,x)}\operatorname{clip}\left[3b_{i}\left|\frac{\Delta(s,x)}{2H}\right.\right]$

### v2

@@ 1 Introduction
Notations. We will consistently use capital letters ($V$) to denote random variables, lowercase letters ($v$) for their values, and cursive $\mathcal{V}$ to denote their domains. Fix indices $i,j\in\mathbb{N}$. We use bold capital letters ($\boldsymbol{V}$) to denote a set of random variables and let $|\boldsymbol{V}|$ denote its cardinality of set $\boldsymbol{V}$. Finally, $\boldsymbol{1}_{\boldsymbol{Z}=\boldsymbol{z}}$ is an indicator function that returns $1$ if event $\boldsymbol{Z}=\boldsymbol{z}$ holds true; otherwise, it returns 0.
@@ 2.1 Potential-Based Reward Shaping
where $r_{h}$ is the reward function in $\mathcal{M}$; $\phi_{h}(\cdot):\mathcal{S}\mapsto\mathbb{R}$ is a real valued potential function and $\phi_{H}(s)=0$. Then every optimal policy in $\mathcal{M}$ will also be an optimal policy in $\mathcal{M}^{\prime}$, and vice versa.
@@ 3 Robust Potential Functions from Confounded Offline Data
where $\widetilde{\mathcal{R}}_{h}(s,x)=\mathbb{E}[Y_{h}|S_{h}=s,X_{h}=x]$, $\widetilde{\mathcal{T}}_{h}(s,x,s^{\prime})=P_{h}(S_{h+1}=s^{\prime}|S_{h}=s,X_{h}=x)$ and $P_{h}(x|s)=P_{h}(X_{h}=x|S_{h}=s)$ are all empirical estimations from the offline dataset. $b$ is a known upper bound on the reward signal, $Y_{h}\leq b$. Similar to the bounding approach developed in (Zhang & Bareinboim, 2024), we apply the above bounds to the Bellman Optimal Equation (Sutton & Barto, 2018) and arrive at the following Causal Bellman Optimal Equation that calculates the optimal state value optimistically from the confounded offline datasets.
@@ 4 Efficient Online Learning with PBRS
where we use $\hat{\mathbb{P}}_{h}^{k_{i}}$ to denote the empirical transition of episode $k_{i}$. That is, for the value of a function mapping from state space to real numbers, $f:\mathcal{S}\mapsto\mathbb{R}$,its value given the next state $s^{\prime}$ in episode $k_{i}$ as input is denoted as, $f(s^{\prime})=(\hat{\mathbb{P}}_{h}^{k_{i}}f)(s,x)$. Similarly, we write the expected value of such function $f$ over the transitions of an online CMDP under reward shaping, $\mathcal{M}^{\prime}$,as $({\mathbb{P}}_{h}f)(s,x)=\mathbb{E}_{\mathcal{M}^{\prime}_{\pi}}[f(s^{\prime})|s,x]$. $\alpha_{t}^{i}$ is the cumulative learning rates defined as $\alpha_{t}^{0}=\Pi_{j=1}^{t}(1-\alpha_{j})$ and $\alpha_{t}^{i}=\alpha_{i}\Pi_{j=i+1}^{t}(1-\alpha_{j})$. Note that we also assume deterministic reward functions for notation wise simplicity (Jin et al., 2018). With the notations above, we bound the difference between the learned Q-value and the optimal Q-value with high probability in both direction.
@@
With probability at least 12p1-2p, the difference between the learned Q-value at the beginning of episode $k$ and step $h$ and the optimal Q-value can be bounded as $0\leq Q^{k}_{h}(s,x)-Q^{*}_{h}(s,x)\leq\alpha_{t}^{0}(-Q^{*}_{h}(s,x))+\sum_{i=1}^{t}\alpha^{i}_{t}\left[(\hat{\mathbb{P}}_{h}^{k_{i}}V_{h+1}^{k_{i}}-\hat{\mathbb{P}}_{h}^{k_{i}}V_{h+1}^{*})(s,x)\right]+3b_{t}$ where $b_{t}$ is the UCB bonus as in Algo. 1.
@@
[…]
@@ 5 Experiments
We use tabular state-action representations and the rewards are all deterministic in those environments. There is a step penalty of 0.1-0.1, +0.2+0.2 for getting a coin, 0 for reaching the goal, and 1-1 for touching the lava. The episode ends immediately when either a goal/lava grid is reached or the episode length hits the horizon limit. In windy grid worlds, the state transition is determined by the resultant force of both agent’s action and the wind direction if the agent tries to move. For example, if the agent wants to move right when there is a north wind (wind blowing from the north), the agent will end up being in the lower right grid concerning its original location instead of the right-hand side grid. However, the wind direction can only be observed by more capable behavioral agents. In our collected offline datasets, wind direction is thus a hidden confounder forming a CMDP. See App. D for detailed experiment setups and more results.
@@ Acknowledgements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here. One major reason is that the potential-based reward-shaping framework we considered does not alter the optimal policy of the agent, which is an inherent property of the problem. Our proposed method solely serves as an acceleration for training such agents.
@@ D.1 Experiment Setups
Regarding the environment parameters, for Windy Empty World, the episode length is set to $15$ while the LavaCross series has a horizon of $20$. To compensate for the hard exploration situation, we allow random initial starting states over the whole map walkable area. For training steps, we set a total of 100100K environment steps for Windy Empty World and 2020K for the LavaCross series. We represent the wind distribution in those windy Grid World as a five-tuple meaning the probability mass for (west wind, north wind, east wind, south wind, no wind). We list them as follows,
  1. Windy Empty World: $(.1,.1,.1,.1,.6)$
  2. Windy LavaCross (easy): $(0,.8,0,0,.2)$
  3. Windy Lava Cross (hard): $(0,.7,0,0,.3)$
[…]
@@ Appendix G Useful Results for Proofs
[…]
@@ H.2 Causal Bellman Optimal Equation Proof
where $\widetilde{\mathcal{R}}_{h}(s,x)=\mathbb{E}[Y_{h}|S_{h}=s,X_{h}=x]$, $\widetilde{\mathcal{T}}_{h}$ is shorthand for $\widetilde{\mathcal{T}}_{h}(s,x,s^{\prime})=P(S_{h+1}=s^{\prime}|S_{h}=s,X_{h}=x)$ and $P(x|s)=P_{h}(X_{h}=x|S_{h}=s)$ are estimated from the offline dataset. And $b$ is a known upper bound on the reward signal, $Y_{h}\leq b$. In this step, we upper bound the next state transition by assuming the best case that for the action not taken with probability $P_{h}(\neg x|s)$, the agent transits with probability $1$ the best possible next state, $\max_{s^{\prime\prime}}V^{*}_{h+1}(s^{\prime\prime})$.
@@
Let the initial difference between the value vectors be $\overline{V}_{1}-\overline{V}_{2}||_{\infty}^{w}\leq c$ where $c>0$ is a constant. Then we have, for any state $s\in\mathcal{S}_{k(s)}$ where $k(s)$ means the index of the state set that $s$ belongs to,
@@
[…]
for all $\overline{V}_{1},\overline{V}_{2}$ satisfying $\overline{V}_{1}-\overline{V}_{2}||\leq c$. Thus, $T$ is a contraction mapping with respect to a weighted max norm.
@@ H.3 Q-UCB with Shaping Regret Analysis Details
where $\mathbb{P}_{h}$ denotes the expected transition w.r.t the true confounded MDP and $\hat{\mathbb{P}}_{h}^{k}$ denotes the sample transition experienced by the agent at episode $k$ step $h$.
@@
By Lem. H.3, we have that with probability at least 12p1-2p for all $(s,x,h,k)\in\mathcal{S}\times\mathcal{X}\times[H]\times[K]$,
@@
    By Lem. H.3 and together with the discussion above for the other side, we know $(\hat{\mathbb{P}}_{h}^{k_{i}}\phi-{\mathbb{P}}_{h}\phi)(s,x)\geq-\frac{b_{t}}{2}$ with high probability. While by Lem. G.1, $\sum_{i=1}^{t}\alpha_{t}^{i}b_{i}\geq b_{t}$, thus, we have,
@@
We can split trajectories by whether the high probability event in Lem. H.4 happens. When Lem. H.4 happens with at least 12p1-2p probability, Lem. H.7 is also satisfied and we have,
@@
[…]
@@
[…]